how do I determine the x and y intercepts of the graph of x=3y^2-9 ??
Thanks

Answer by venugopalramana(3286)

how do I determine the x and y intercepts of the graph of x=3y^2-9 ??
x intercept means where the curve will cut x axis and the distance of that point of intersection from origin or x coordinate.when the curve cuts the x axis obviously that point of intersection will have zero y coordinate .hence we can find the x intercept by putting y=0 in the given equation and solving for x...............similarly
y intercept means where the curve will cut y axis and the distance of that point of intersection from origin or y coordinate.when the curve cuts the y axis obviously that point of intersection will have zero x coordinate .hence we can find the y intercept by putting x=0 in the given equation and solving for y...............so
x=3y^2-9 =0
3y^2=9
y^2=9/3=3
y=+sq.rt3 and -sqrt3..we have 2 y intercepts that is the curve cuts y axis at 2 places namely +sq.rt3 and -sqrt3..
dimilarly putting y=0 we get
x=-9...so there is one place of intersection with x axis and the xintercept
is -9